OpenStudy (anonymous):

A pole that is 15 feet tall casts a shadow 33 feet long. At the same time, another pole casts a shadow having a length of 22 feet. What is the height, in feet, of the second pole

OpenStudy (anonymous):

 The relationship between the two triangles is 15/33 = x/22 multiply 22 on both sides to get 330/33 = x x=10 feet

OpenStudy (mertsj):